OpenStudy (anonymous):

Find the remainder when f(x) is divided by (x - k) f(x) = 3x3 - 4x2 - 3x + 14; k= 3

OpenStudy (campbell_st):

you can use the remainder theorem and find f(3) with will be the remainder for dividing f(x) by ( x - 3)

OpenStudy (anonymous):

so do you know which one is the answer? 50 68 -12 112

OpenStudy (nurali):

f(x) = 3x3 - 4x2 - 3x + 14 f(3)=3(3)^3-4(3)^2-3(3)+14 f(3)=81-36-9+14 f(3)=50